OpenStudy (anonymous):

The table shows how the weight of a baby horse is increasing over time. Age (in days) Weight (in pounds) 2 116 4 124 6 132 8 140 10 148 According to the table, what was the baby horse's weight at an age of 5 days? A. 127 pounds B. 128 pounds C. 130 pounds D. 126 pounds

OpenStudy (godlovesme):

can u find the slope

OpenStudy (godlovesme):

\[y2-y1/x2-x1\]

OpenStudy (anonymous):

ok I have a question, when it asks for that does it want like half of it? so their 8 lbs apart. so half of that would be 126 lbs? Am I on the right track with it??

OpenStudy (anonymous):

Agree with @Godlovesme, to find it at 5 years you have to find the rate of growth

OpenStudy (godlovesme):

i should look like this 124-116/ 4-2 ---> 8/2 --> 4 u forgot to divide it by 2 @Tallan

OpenStudy (anonymous):

124 - 116/4-2 = 8/2 = 4 132-124/6-4 = 8/2 = 4 140-132/8-6 = 8/2 = 4 148-140/10-8 = 8/2 = 4

OpenStudy (godlovesme):

wow nice job!! :) @BPDlkeme234

OpenStudy (anonymous):

?I'm there!! I'm totally there!

OpenStudy (anonymous):

continue @Godlovesme

OpenStudy (godlovesme):

loll ;D @Tallan got it?

OpenStudy (anonymous):

now we know we have a linear model!

OpenStudy (anonymous):

So it would be B right?

OpenStudy (godlovesme):

yep :)

OpenStudy (anonymous):

and we also know the slope!

OpenStudy (anonymous):

m = 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hey, I am not taking this away from @Godlovesme

OpenStudy (godlovesme):

lol i gave u the medal :) @BPDlkeme234

OpenStudy (anonymous):

are you sure? @Godlovesme , I can leave it with you or provide hint!

OpenStudy (anonymous):

hint y = 4x +108

OpenStudy (godlovesme):

yes and god job setting up the actual eqn

OpenStudy (anonymous):

tnx @Godlovesme

OpenStudy (godlovesme):

np

OpenStudy (anonymous):

so, the person still needs the answer, i.e. substitute x=5, take it away @Godlovesme

OpenStudy (anonymous):

The answer would be B 128 lbs

OpenStudy (godlovesme):

he/she already got the answer though @BPDlkeme234

OpenStudy (anonymous):

cool! great job @tallan and @Godlovesme

OpenStudy (anonymous):

im a he :P @Godlovesme

OpenStudy (godlovesme):

thx for the medal guys ^.^ lol @Tallan xP

OpenStudy (anonymous):

P.S. Derivation of the equation was simple: y-y1 = m(x-x1), you already gave me the slope!

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

Tallan has a sense of humor, I like that!

OpenStudy (anonymous):

I try! xD

OpenStudy (godlovesme):

;D lolz i'm leaving dont forget to close the question :p @Tallan

OpenStudy (anonymous):

Alright take care @Godlovesme :)

OpenStudy (godlovesme):

ight :D @Tallan